solve for x
ax-x=c
Here is what I've come up with... x=c/a-1
Is this correct?

Answer by Maths68(1474)

You are right well done!
Keep it up!!
solve for x
ax-x=c
x(a-1)=c
x(a-1)/(a-1)=c/(a-1)
cancel out (a-1) from left hand side of the equation
x=c/(a-1)